OpenStudy (sphott51):

write the equation for the nth term of the geometric sequence 12, -3, 3/4, ...

Directrix (directrix):

The common ratio is -4. The sequence will continue as: 12, -48, 192, -768, 3072, -12288, 49152 ... The terms alternate in sign. Therefore, the general term will have a (-4)^n in it. If n = 1, then (-4)^n becomes (-4)^1 = -4 If n = 2, then (-4)^n becomes (-4)^2 = 16 Looking at the first two terms: 12, -48, -3*(-4) = 12 -3*(16) = -48 Test others of the terms. The nth term looks to be: -3 * (-4)^n See what you think. @SPHott51